Find the length of a diagonal of a square if its area is 162 square feet
Found 2 solutions by Boreal, addingup:
Answer by Boreal(15235)

For the diagonal, one needs the side length, and then multiply that by sqrt(2).
s^2=162
s=sqrt(162)=sqrt(81)*sqrt(2)=9sqrt(2) feet.
s sqrt(2)=9 sqrt(2)*sqrt(2)=18 feet.

Answer by addingup(3677)

√162 = 12.728
````````````````````````
√12.728^2+12.728^2 = diagonal
√324 = diagonal Use your calculator.